Suppose an atom of Osmium at rest emits an X-ray photon of energy 6.8 keV. Calculate the “recoil” momentum and kinetic energy of the atom.
1
Expert's answer
2020-06-15T09:49:02-0400

As,system has initial momentum is zero and no external interaction there,Momentum of the system is conserved.

Therefore, momentum

"p_{os}+p_{\\gamma}=0\n\n\n\u200b"

We have given that, emitted photon has energy is "8.6keV"

Since,

"E_{\gamma}=p_{\gamma}c\\ \implies p_{\gamma}=\frac{E_{\gamma}}{c}=8.6keV/c"


Thus,

"p_{os}=-8.6keV\/c"

Kinetic energy is

"K=\frac{p^2}{2m}"


Hence,

"K_{os}=2.088\times 10^{-4}eV"